# Photon pair generation via down-conversion in III-V semiconductor microrings: modal dispersion and quasi-phase matching

### v1

@@ II.5.2 Lowest-order squeezing regime
where within the approximations identified above the pair generation operator is the same as previously defined in Eq. (64), and the norm of $\beta$ is still given by Eq. (75), but now $\beta$ (the “squeezing parameter” [21]), no longer immediately gives the probability of generating pairs, and $\varphi^{\mathrm{II}}_{\lambda\lambda^{\prime}}(k_{1},k_{2})$ is more properly identified as the “joint spectral amplitude.” The ket for the generated light in the LOS regime is given by a multi-mode squeezed vacuum state
@@
In order to calculate these moments we perform a Schmidt decomposition [20, 19] of the joint spectral amplitude in Eq. (87). This is achieved by decomposing the matrix $\bm{J}$ as
@@
An expression similar to Eq. (59) for the generation efficiency in the CW limit, and similar to Eq. (78) for the efficiency for the pair-regime, can then be introduced here for the LOS regime,
@@
As we will see in our sample calculation, $\mathcal{E}^{SI}_{\mathrm{pulse}}$ coincides with $\mathcal{E}^{SI}_{\mathrm{pulse;pair-regime}}$ for sufficiently low pulse energies and powers, as we would expect. To study the separability of the squeezed state in Eq. (87) we calculate its Schmidt number [20, 19]
@@
where $r_{n}$ are the entries of the diagonal matrix $\bm{R}$ in Eq. (92). Within the approximations of this subsection, the Schmidt number in the LOS regime is unchanged from that in the pair-regime, since the same pair generation operator from Eq. (64) appears in Eq. (63) from the pair-regime and expression (81) from the LOS regime for the output ket. That is, within the approximations made here the joint spectral amplitude in the LOS regime is the same as the biphoton wave function in the pair-regime.
@@ III Sample Calculations
Figure 3: (a) Electric field $x$ component of the fundamental TE mode ($1550$ nm) for signal and idler, and (b) Electric field $z$ component of the fundamental TM ($775$ nm) mode for the pump. Simulation done on Lumerical for InGaP with a height of $102$ nm, width of $1300$ nm and bend radius of $20\,\mu$m; the centre of the ring is at negative $x$. The plotted fields are normalized by the maximum of the modulus squared of the total electric field.
Consider first CW excitation at a frequency within the ring resonance at $\omega_{P}$, and the generation of signal and idler fields within the ring resonances centered at $\omega_{S}$ and $\omega_{I}$ respectively. This is possible only if the quasi-phase matching condition from Eq. (41) is satisfied, which can be written as
@@
As a first estimate of ideal operating points, we put $\omega_{S}\approx\omega_{I}\approx\frac{1}{2}\omega_{P}$, and neglect the quasi-phase matching term in Eq. (99), giving $\Delta\kappa=0$, or
@@
In the calculations below we adopt realistic quality factors: intrinsic quality factors of $1\times 10^{6}$ at the signal and idler frequencies, and of $1\times 10^{5}$ at the pump frequency [7, 11, 30]. We consider a range of couplings, but at critical coupling the simulated rings (with $\mathcal{R}=30\,\mu$m) have a finesse of around $\mathcal{F}_{P}^{\mathrm{ac}}\approx 50$ and $\mathcal{F}_{S}^{\mathrm{ac}}=\mathcal{F}_{I}^{\mathrm{ac}}\approx 1300$ (where $\mathcal{F}_{u}^{\lambda}={v_{u}^{\lambda}}/{2\bar{\Gamma}_{u}\mathcal{R}}$ [20]): the field intensity in the ring is amplified by a factor of $\sim 15$ for the pump frequency compared to the incident field, and the enhancement factor for the signal and idler is $\sim 425$ [20].
In the following subsections, we present results for the efficiencies for both CW and pulsed excitation.
@@ III.1 CW excitation
We now go beyond the simple estimate of Eq. (101) in determining for what structures a large pair production rate can be achieved, and how large that rate can be. We consider structures close to those identified by the intersections in Fig. 4, and of course in evaluating the rate (Eq. (52)) need only consider signal and idler resonances for which the quasi-phase matching condition from Eq. (42) is satisfied. Each resulting $R^{SI}_{\lambda\lambda^{\prime}}$ is optimized when
@@
The structures yielding high generation efficiencies are given by the darker spots in Fig. 5 (a-f), which we refer to as “hotspots,” and at these in general only one set of quasi-phase matched ring resonances contributes significantly to the total rate. The highest efficiencies from each family of structures investigated in Fig. 5 are tabulated in Tab. 1. For excitation at wavelength $\lambda_{0}\approx 775$ nm, the optimized generation efficiencies range from $0.53-1.47\times 10^{-5}$. These efficiencies are in accord with recent experimental results [8].
@@
We note that having to satisfy one of the quasi-phase matching (QPM) conditions $\Delta\kappa\pm 2/\mathcal{R}=0$ instead of the usual phase matching (UPM) condition $\Delta\kappa=0$ (see Eqs. (100)-(101) and the discussion following) does not necessarily lead to an increase in the generation efficiency for a given family of structures of the type considered here.
@@
We give a quantitative example: for the InGaP ring with radius of 40 microns ((f) and (g) from Fig. 5), the single UPM hotspot is a structure with a height of $102$ nm, a width of $1206$ nm, and an efficiency of $0.16\times 10^{-5}$; the minus and plus QPM hotspots for the same $102$ nm height have widths of $1142$ nm and $1262$ nm respectively, and have efficiencies of $0.81\times 10^{-5}$ and $0.06\times 10^{-5}$ respectively. The mode numbers of the resonances contributing the most to the rates for the UPM condition are $m_{P}=610$, $m_{S}=309$, and $m_{I}=301$ (with wavelengths $\lambda_{P}=774.88$ nm, $\lambda_{S}=1536.20$ nm, and $\lambda_{I}=1563.56$ nm). The mode numbers of the resonances contributing the most to the rates for the minus and plus QPM hotspots are, respectively: $m_{P}=609$, $m_{S}=315$, and $m_{I}=292$ (with wavelengths $\lambda_{P}=774.97$ nm, $\lambda_{S}=1511.64$ nm, and $\lambda_{I}=1590.24$ nm), and $m_{P}=610$, $m_{S}=318$, $m_{I}=294$ (with wavelengths $\lambda_{P}=775.30$ nm, $\lambda_{S}=1510.51$ nm, and $\lambda_{I}=1592.85$ nm).
@@
  AlGaAs a) 20 $938\times 105$ $1.47\times 10^{-5}$ $57.2$ $0.543$ [--] $170$
b) 30 $986\times 104$ $0.93\times 10^{-5}$ $36.2$ $0.670$ [$+$] $-61$
c) 40 $958\times 105$ $0.82\times 10^{-5}$ $32.0$ $0.563$ [--] $-74$
InGaP d) 20 $1266\times 104$ $1.58\times 10^{-5}$ $61.7$ $0.556$ [--] $109$
e) 30 $1142\times 100$ $0.53\times 10^{-5}$ $20.8$ $0.734$ [$+$] $-157$
f) 40 $1142\times 102$ $0.81\times 10^{-5}$ $31.5$ $0.578$ [--] $-51$
Table 1: Largest generation efficiencies of the different families of structures in the subfigures of Fig. 5, with the respective effective areas and the values of $\Delta\omega$ for the set of resonances that contributes the most to the rates. The $\pm$ in square brackets indicates which quasi-phase matching condition is met for the resonances that contribute the most to the generation rate. We have taken $\bar{n}_{P}=\bar{n}_{S}=\bar{n}_{I}=\bar{n}=3.5$, and $\bar{v}_{P}=\bar{v}_{S}=\bar{v}_{I}=c/\bar{n}$.
Figure 7: Spectrum of the signal photons leaving the actual waveguide for the optimized AlGaAs microring with $\mathcal{R}=30\,\mu$m, cross-sectional dimensions of $986\,\mathrm{nm}\times 104$ nm (b) from Tab. 1). Critical coupling is assumed, $\bar{\chi}^{(2)}=220$ pm/V, $\lambda_{S}\in[1500\,\mathrm{nm},2\times\lambda_{P}]$ nm, $\lambda_{P}=774.82$ nm, $\delta\omega=0$, $Q^{\mathrm{(load)}}_{S}=Q^{\mathrm{(load)}}_{I}=5\times 10^{5}$, $Q^{\mathrm{(load)}}_{P}=5\times 10^{4}$. The dashed lines represent the centre frequencies of the different signal resonance bins $\omega_{S}$. The idler frequency is given by $\omega_{2}=\omega_{0}-\omega_{1}$, the pump mode number is $m_{P}=474$, and the idler mode numbers are given by $m_{I}=m_{P}-m_{S}\pm 2$.
@@
Here we have used Eq. (53) and Eq. (118) for $R^{SI}_{\mathrm{ac,ac}}$ and $\hat{R}^{SS}_{\mathrm{ac,ac}}$ respectively, changed variables, and integrated over the idler frequency $\omega_{Ik_{2}}^{\mathrm{ac}}$ with the Dirac delta function. Thus $f_{\mathrm{ac,ac}}(\omega_{1})$ can be identified as the spectrum of the signal photon when both photons exit the actual channel, and in Fig. 7 we show it for the optimized AlGaAs ring with $\mathcal{R}=30\,\mu$m, restricting ourselves to three resonances, the centre one being the one that contributes the most to the rate; the other resonances contribute very little to the rate. We see that the spectrum is highly peaked for one resonance, which dominates the rate, and this is true in general for the hotspots we see in Fig. 5. The double peaks in the spectrum come from the two field enhancement factors that appear in the spectrum $f_{\mathrm{ac,ac}}(\omega_{1})$: these two field enhancement factors are Lorentzians peaked at $\omega_{1}=\omega_{S}$ and $\omega_{1}=\omega_{0}-\omega_{I}$ respectively. The first of these corresponds to the signal being on resonance, and the second to the idler being on resonance, $\omega_{0}-\omega_{1}=\omega_{I}$. If the frequency bin matching condition were met exactly, i.e. $\omega_{S}=\omega_{0}-\omega_{I}$ and $\Delta\omega=0$, the two Lorentzians would peak at the same frequency $\omega_{1}$, and only a single peak per resonance would appear in the spectrum; usually this does not occur, and two peaks are typically associated with each resonance, although in some instances they are not resolved.
@@
We stress that the rate equation and conversion efficiencies discussed thus far are in the weak pumping limit; at large enough pump powers the Fermi’s Golden Rule calculation presented here will break down. This is more easily explored, and may be more relevant, in the scenario of pump pulse excitation, which we consider in the next subsection.
@@ III.2 Pulsed excitation
Figure 8: Joint spectral amplitudes (we plot $\big{(}|\varphi^{\mathrm{II}}(k_{1},k_{2})|^{2}/\mathrm{max}\left[|\varphi^{\mathrm{II}}(k_{1},k_{2})|^{2}\right]\big{)}$) from a ring of radius $\mathcal{R}=30\,\mu$m, using resonances $\lambda_{P}=774.82\,\mathrm{nm}$, $\lambda_{S}=1524.57\,\mathrm{nm}$, $\lambda_{I}=1575.54\,\mathrm{nm}$, and with $\Delta\omega/2\pi=-61$ MHz and $\delta\omega=0$. The pump pulses are Gaussian with various FWHM $\tau_{P}$ as indicated, and peak powers of $P_{P}=10\,\mu$W; they have energies $E_{P}=$ (a) $0.16$ fJ, (b) $1.6$ fJ, (c) $16$ fJ, (d) $0.16$ pJ. The Schmidt numbers of the resulting joint spectral amplitudes are $K=$ (a) $1.00044$, (b) $1.0036$, (c) $1.19$, (d) $3.92$. The total number of signal photons in the outgoing states are $N_{SI}=$ (a) $5.61\times 10^{-4}$, (b) $0.0535$, (c) $3.29$, (d) $62.94$, and $|\beta|^{2}=$ (a) $5.61\times 10^{-4}$, (b) $0.0526$, (c) $2.00$, (d) $23.02$.
Due to the similarities of the equations for CW and pulsed excitation, we can expect that the optimized structures yielding high efficiencies for CW excitation also yield high efficiencies for pulsed excitation. So we focus on one of our optimized rings from Tab. 1: the AlGaAs ring of radius $\mathcal{R}=30\,\mu$m and cross-sectional dimensions of $986\,\mathrm{nm}\times 104$ nm. We take the coupling to be critical, and consider Gaussian pulses,
@@
with temporal full width half maximum values $\tau_{P}$ ranging from $15$ ps to $15$ ns, always with a peak power of $10$ $\mu$W; the pump pulse energies then range from $0.16$ fJ to $0.16$ pJ. The set of resonances identified by the mode numbers $m_{P}=474$, $m_{S}=244$ and $m_{I}=232$, with centre wavelengths $\lambda_{P}=774.82\,\mathrm{nm}$, $\lambda_{S}=1524.57\,\mathrm{nm}$, and $\lambda_{I}=1575.54\,\mathrm{nm}$, contributes most to the generation of photon pairs. Here the plus quasi-phase matching condition is satisfied, and $\Delta\omega/2\pi=-61$ MHz: this is comparable to the sum of the linewidths of the signal and idler $(\bar{\Gamma}_{S}+\bar{\Gamma}_{I})/2\pi=387\,\mathrm{MHz}$.
The joint spectral amplitudes are shown in Fig. 8. We see that for realistic pump pulses such a structure can be used to generate pairs that have both low ($1.00044$), and high ($3.92$) Schmidt numbers, depending on the duration of the pump pulse. By varying the coupling even higher and lower Schmidt numbers can be obtained. For the shorter and lower energy pulses (Fig. 8 (a) and (b)), the number of generated signal photons $N_{SI}$ calculated from Eq. (96) essentially equals $|\beta|^{2}$, as expected in the pair-regime from Eq. (75). But for longer and higher energy pulses (Fig. 8 (c) and (d)) we see that the approximation of the pair-regime fails, as $N_{SI}$ becomes significantly larger than $|\beta|^{2}$.
Figure 9: Scaling of the generation efficiency for varying pulse energies for the optimized AlGaAs ring of radius $\mathcal{R}=30\,\mu$m, where $\lambda_{P}=774.82\,\mathrm{nm}$, $\lambda_{S}=1524.57\,\mathrm{nm}$, $\lambda_{I}=1575.54\,\mathrm{nm}$ and $\Delta\omega/2\pi=-61$ MHz. The pump pulse duration is fixed: $\tau_{P}=1.5$ ns. The LOS approximation (solid red line) diverges from the approximation from the pair-regime (dashed black line) for pulse energies of around 1–10 fJ, corresponding to peak powers of 0.63–6.3 $\mu$W.
To investigate this further, with the same optimized 30 $\mu$m AlGaAs ring from Tab. 1 we consider excitation by pump pulses all of duration $\tau_{P}=1.5$ ns, but with varying pulse energies, (and hence peak powers). In Fig. 9 we plot the predicted efficiencies from our approximation for both the pair-regime and LOS regime, and we see that for relatively low and realistic pulse energies not only does the pair-regime approximation fails, but the LOS approximation predicts efficiencies so large that pump depletion – which has been neglected here – would have to be taken into account to determine a reliable estimate. Future work will include both time-ordering corrections, which have been neglected in the simple LOS approximation presented here, and pump depletion.

### v3

@@ II.5.2 Lowest-order squeezing regime
where within the approximations identified above the pair generation operator is the same as previously defined in Eq. (64), and the norm of $\beta$ is still given by Eq. (75), but now $\beta$ (the “squeezing parameter” [23]), no longer immediately gives the probability of generating pairs, and $\varphi^{\mathrm{II}}_{\lambda\lambda^{\prime}}(k_{1},k_{2})$ is more properly identified as the “joint spectral amplitude.” The ket for the generated light in the LOS regime is given by a multimode squeezed vacuum state
@@
To calculate these moments we perform a Schmidt decomposition [22, 21] of the joint spectral amplitude in Eq. (87). This is achieved by decomposing the matrix $\bm{J}$ as
@@
An expression similar to Eq. (59) for the generation efficiency in the cw limit, and similar to Eq. (78) for the efficiency for the pair regime, can then be introduced here for the LOS regime,
@@
As we will see in our sample calculation, $\mathcal{E}^{SI}_{\mathrm{pulse}}$ coincides with $\mathcal{E}^{SI}_{\mathrm{pulse;pairregime}}$ for sufficiently low pulse energies and powers, as we would expect. To study the separability of the squeezed state in Eq. (87) we calculate its Schmidt number [22, 21]
@@
where $r_{n}$ are the entries of the diagonal matrix $\bm{R}$ in Eq. (92). Within the approximations of this section, the Schmidt number in the LOS regime is unchanged from that in the pair regime, since the same pair generation operator from Eq. (64) appears in Eq. (63) from the pair regime and expression (81) from the LOS regime for the output ket. That is, within the approximations made here the joint spectral amplitude in the LOS regime is the same as the biphoton wave function in the pair regime.
@@ III Sample Calculations
Figure 3: (a) Electric-field $x$ component of the fundamental TE mode ($1550$ nm) for signal and idler, and (b) Electric-field $z$ component of the fundamental TM ($775$ nm) mode for the pump. Simulation done on Lumerical for InGaP with a height of $102$ nm, width of $1300$ nm and bend radius of $20\,\mu$m; the centre of the ring is at negative $x$. The plotted fields are normalized by the maximum of the modulus squared of the total electric field.
Consider first cw excitation at a frequency within the ring resonance at $\omega_{P}$, and the generation of signal and idler fields within the ring resonances centered at $\omega_{S}$ and $\omega_{I}$ respectively. This is possible only if the quasi-phase-matching condition from Eq. (41) is satisfied, which can be written as
@@
As a first estimate of ideal operating points, we put $\omega_{S}\approx\omega_{I}\approx\frac{1}{2}\omega_{P}$, and neglect the quasi-phase-matching term in Eq. (99), giving $\Delta\kappa=0$, or
@@
In the calculations below we adopt realistic quality factors: intrinsic quality factors of $1\times 10^{6}$ at the signal and idler frequencies, and of $1\times 10^{5}$ at the pump frequency [8, 12, 31]. We consider a range of couplings, but at critical coupling the simulated rings (with $\mathcal{R}=30\,\mu$m) have a finesse of around $\mathcal{F}_{P}^{\mathrm{ac}}\approx 50$ and $\mathcal{F}_{S}^{\mathrm{ac}}=\mathcal{F}_{I}^{\mathrm{ac}}\approx 1300$ (where $\mathcal{F}_{u}^{\lambda}={v_{u}^{\lambda}}/{2\bar{\Gamma}_{u}\mathcal{R}}$ [22]): the field intensity in the ring is amplified by a factor of $\sim 15$ for the pump frequency compared with the incident field, and the enhancement factor for the signal and idler is $\sim 425$ [22].
In the following sections, we present results for the efficiencies for both cw and pulsed excitation.
@@
We now go beyond the simple estimate of Eq. (101) in determining for what structures a large pair production rate can be achieved, and how large that rate can be. We consider structures close to those identified by the intersections in Fig. 4, and of course in evaluating the rate (Eq. (52)) need only consider signal and idler resonances for which the quasi-phase-matching condition from Eq. (42) is satisfied. Each resulting $R^{SI}_{\lambda\lambda^{\prime}}$ is optimized when
@@
The structures yielding high generation efficiencies are given by the red spots in Fig. 5 (a-f), which we refer to as “hotspots,” and at these in general only one set of quasi-phase-matched ring resonances contributes significantly to the total rate. The highest efficiencies from each family of structures investigated in Fig. 5 are tabulated in Tab. 1. For excitation at wavelength $\lambda_{0}\approx 775$ nm, the optimized generation efficiencies range from $0.67-1.39\times 10^{-5}$. These efficiencies are in accord with recent experimental results [7].
@@
We note that having to satisfy one of the quasi-phase-matching (QPM) conditions $\Delta\kappa\pm 2/\mathcal{R}=0$ instead of the usual phase matching (UPM) condition $\Delta\kappa=0$ (see Eqs. (100)-(101) and the discussion following) does not necessarily lead to an increase in the generation efficiency for a given family of structures of the type considered here.
@@
We give a quantitative example: for the InGaP ring with radius of 40 microns ((f) and (g) from Fig. 5), the single UPM hotspot is a structure with a height of $102$ nm, a width of $1278$ nm, and an efficiency of $0.28\times 10^{-6}$; the minus and plus QPM hotspots for the same $102$ nm height have widths of $1206$ nm and $1358$ nm respectively, and have efficiencies of $7.72\times 10^{-6}$ and $0.39\times 10^{-6}$ respectively. The mode numbers of the resonances contributing the most to the rates for the UPM condition are $m_{P}=612$, $m_{S}=314$, and $m_{I}=298$ (with wavelengths $\lambda_{P}=774.87$ nm, $\lambda_{S}=1522.93$ nm, and $\lambda_{I}=1577.52$ nm). The mode numbers of the resonances contributing the most to the rates for the minus and plus QPM hotspots are, respectively: $m_{P}=611$, $m_{S}=315$, and $m_{I}=294$ (with wavelengths $\lambda_{P}=774.96$ nm, $\lambda_{S}=1514.96$ nm, and $\lambda_{I}=1586.52$ nm), and $m_{P}=613$, $m_{S}=319$, $m_{I}=296$ (with wavelengths $\lambda_{P}=774.78$ nm, $\lambda_{S}=1511.24$ nm, and $\lambda_{I}=1589.86$ nm).
@@
  AlGaAs a) 20 $1026\times 106$ $1.04\times 10^{-5}$ $40.5$ $0.519$ [--] $-340$
b) 30 $894\times 103$ $0.67\times 10^{-5}$ $26.3$ $0.563$ [--] $-324$
c) 40 $1186\times 108$ $0.88\times 10^{-5}$ $34.4$ $0.537$ [--] $-27$
InGaP d) 20 $1386\times 104$ $1.39\times 10^{-5}$ $54.4$ $0.568$ [--] $-173$
e) 30 $1194\times 102$ $1.17\times 10^{-5}$ $45.6$ $0.557$ [--] $-17$
f) 40 $1206\times 102$ $0.77\times 10^{-5}$ $30.1$ $0.567$ [--] $113$
Table 1: Largest generation efficiencies of the different families of structures in the subfigures of Fig. 5, with the respective effective areas and the values of $\Delta\omega$ for the set of resonances that contributes the most to the rates. The $\pm$ in square brackets indicates which quasi-phase-matching condition is met for the resonances that contribute the most to the generation rate. We have taken $\bar{n}_{P}=\bar{n}_{S}=\bar{n}_{I}=\bar{n}=3.5$, and $\bar{v}_{P}=\bar{v}_{S}=\bar{v}_{I}=c/\bar{n}$.
Figure 7: Spectrum of the signal photons exiting the actual waveguide for the optimized AlGaAs microring with $\mathcal{R}=30\,\mu$m, cross-sectional dimensions of $894\,\mathrm{nm}\times 103$ nm (b) from Tab. 1). Critical coupling is assumed, $\bar{\chi}^{(2)}=220$ pm/V, $\lambda_{S}\in[1500\,\mathrm{nm},2\times\lambda_{P}]$ nm, $\lambda_{P}=775.23$ nm, $\delta\omega=0$, $Q^{\mathrm{(load)}}_{S}=Q^{\mathrm{(load)}}_{I}=5\times 10^{5}$, $Q^{\mathrm{(load)}}_{P}=5\times 10^{4}$. The dashed lines represent the centre frequencies of the different signal resonance bins $\omega_{S}$. The idler frequency is given by $\omega_{2}=\omega_{0}-\omega_{1}$, the pump mode number is $m_{P}=469$, and the idler mode numbers are given by $m_{I}=m_{P}-m_{S}\pm 2$.
@@
Here we have used Eq. (53) and Eq. (118) for $R^{SI}_{\mathrm{ac,ac}}$ and $\hat{R}^{SS}_{\mathrm{ac,ac}}$ respectively, changed variables, and integrated over the idler frequency $\omega_{Ik_{2}}^{\mathrm{ac}}$ with the Dirac delta function. Thus $f_{\mathrm{ac,ac}}(\omega_{1})$ can be identified as the spectrum of the signal photon when both photons exit the actual channel, and in Fig. 7 we show it for the optimized AlGaAs ring with $\mathcal{R}=30\,\mu$m, restricting ourselves to three resonances, the centre one being the one that contributes the most to the rate; the other resonances contribute very little to the rate. We see that the spectrum is highly peaked for one resonance, which dominates the rate, and this is true in general for the hotspots we see in Fig. 5. The double peaks in the spectrum come from the two field-enhancement factors that appear in the spectrum $f_{\mathrm{ac,ac}}(\omega_{1})$: these two field-enhancement factors are Lorentzians peaked at $\omega_{1}=\omega_{S}$ and $\omega_{1}=\omega_{0}-\omega_{I}$ respectively. The first of these corresponds to the signal being on resonance, and the second to the idler being on resonance, $\omega_{0}-\omega_{1}=\omega_{I}$. If the frequency bin matching condition were met exactly, i.e. $\omega_{S}=\omega_{0}-\omega_{I}$ and $\Delta\omega=0$, the two Lorentzians would peak at the same frequency $\omega_{1}$, and only a single peak per resonance would appear in the spectrum; usually this does not occur, and two peaks are typically associated with each resonance, although in some instances they are not resolved.
@@
We stress that the rate equation and conversion efficiencies discussed thus far are in the weak pumping limit; at large enough pump powers the Fermi’s golden rule calculation presented here will break down. This is more easily explored, and may be more relevant, in the scenario of pump pulse excitation, which we consider in the next section.
@@ III.2 Pulsed excitation
Figure 8: Joint spectral amplitudes (we plot $\big{(}|\varphi^{\mathrm{II}}(k_{1},k_{2})|^{2}/\mathrm{max}\left[|\varphi^{\mathrm{II}}(k_{1},k_{2})|^{2}\right]\big{)}$) from a ring of radius $\mathcal{R}=30\,\mu$m, using resonances $\lambda_{P}=775.23\,\mathrm{nm}$, $\lambda_{S}=1531.64\,\mathrm{nm}$, $\lambda_{I}=1569.75\,\mathrm{nm}$, and with $\Delta\omega/2\pi=-324$ MHz and $\delta\omega=0$. The pump pulses are Gaussian with various FWHM $\tau_{P}$ as indicated, and peak powers of $P_{P}=10\,\mu$W; they have energies $E_{P}=$ (a) $0.16$ fJ, (b) $1.6$ fJ, (c) $16$ fJ, (d) $0.16$ pJ. The Schmidt numbers of the resulting joint spectral amplitudes are $K=$ (a) $1.00031$, (b) $1.0028$, (c) $1.31$, (d) $8.38$. The total number of signal photons in the outgoing states are $N_{SI}=$ (a) $6.50\times 10^{-4}$, (b) $0.0619$, (c) $2.49$, (d) $26.58$, and $|\beta|^{2}=$ (a) $6.50\times 10^{-4}$, (b) $0.0607$, (c) $1.70$, (d) $16.43$.
Due to the similarities of the equations for cw and pulsed excitation, we can expect that the optimized structures yielding high efficiencies for cw excitation also yield high efficiencies for pulsed excitation. So we focus on one of our optimized rings from Tab. 1: the AlGaAs ring of radius $\mathcal{R}=30\,\mu$m and cross-sectional dimensions of $894\,\mathrm{nm}\times 103$ nm. We take the coupling to be critical, and consider Gaussian pulses,
@@
with temporal full width half maximum values $\tau_{P}$ ranging from $15$ ps to $15$ ns, always with a peak power of $10$ $\mu$W; the pump pulse energies then range from $0.16$ fJ to $0.16$ pJ. The set of resonances identified by the mode numbers $m_{P}=469$, $m_{S}=238$ and $m_{I}=229$, with centre wavelengths $\lambda_{P}=775.23\,\mathrm{nm}$, $\lambda_{S}=1531.64\,\mathrm{nm}$, and $\lambda_{I}=1569.75\,\mathrm{nm}$, contributes most to the generation of photon pairs. Here the plus quasi-phase-matching condition is satisfied, and $\Delta\omega/2\pi=-324$ MHz: this is comparable to the sum of the linewidths of the signal and idler $(\bar{\Gamma}_{S}+\bar{\Gamma}_{I})/2\pi=387\,\mathrm{MHz}$.
The joint spectral amplitudes are shown in Fig. 8. We see that for realistic pump pulses such a structure can be used to generate pairs that have both low ($1.00031$), and high ($8.38$) Schmidt numbers, depending on the duration of the pump pulse. By varying the coupling even higher and lower Schmidt numbers can be obtained. For the shorter and lower energy pulses (Fig. 8 (a) and (b)), the number of generated signal photons $N_{SI}$ calculated from Eq. (96) essentially equals $|\beta|^{2}$, as expected in the pair regime from Eq. (75). But for longer and higher energy pulses (Fig. 8 (c) and (d)) we see that the approximation of the pair regime fails, as $N_{SI}$ becomes significantly larger than $|\beta|^{2}$.
Figure 9: Scaling of the generation efficiency for varying pulse energies for the optimized AlGaAs ring of radius $\mathcal{R}=30\,\mu$m, where $\lambda_{P}=775.23\,\mathrm{nm}$, $\lambda_{S}=1531.64\,\mathrm{nm}$, $\lambda_{I}=1569.75\,\mathrm{nm}$ and $\Delta\omega/2\pi=-324$ MHz. The pump pulse duration is fixed: $\tau_{P}=1.5$ ns. The LOS approximation (solid red line) diverges from the approximation from the pair regime (dashed black line) for pulse energies of around 1–10 fJ, corresponding to peak powers of 0.63–6.3 $\mu$W.
To investigate this further, with the same optimized 30 $\mu$m AlGaAs ring from Tab. 1 we consider excitation by pump pulses all of duration $\tau_{P}=1.5$ ns, but with varying pulse energies, (and hence peak powers). In Fig. 9 we plot the predicted efficiencies from our approximation for both the pair regime and LOS regime, and we see that for relatively low and realistic pulse energies not only does the pair regime approximation fails, but the LOS approximation predicts efficiencies so large that pump depletion – which has been neglected here – would have to be taken into account to determine a reliable estimate. Future work will include both time-ordering corrections, which have been neglected in the simple LOS approximation presented here, and pump depletion.